# Optimizing Sequencing Coverage Depth in DNA Storage: Insights From DNA Storage Data

### v1

@@ IV-A Sequencing Coverage Depth in the Noiseless Channel
Theorem 1. For given any channel probability distribution $\boldsymbol{p}_{t}=(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{n}^{(t)})$ and any $K\geq n\geq m\geq 1$, it holds that
@@
$\mathbb{E}[X_{i}]=1-\left(1-p_{i}^{(t)}\right)^{K}\approx 1-e^{K\cdot p_{i}^{(t)}},$
@@
$\mathbb{E}[N]=\mathbb{E}\left[\sum_{i=1}^{n}X_{i}\right]=\sum_{i=1}^{n}\mathbb{E}[X_{i}]=\sum_{i=1}^{n}\left(1-e^{K\cdot p_{i}^{(t)}}\right).$
The value of $Var[N]$: Inspired by [15], we first investigate the number of urns that remain empty after $K$ balls thrown into $n$ urns. Let $N_{0}$ represents the number of urns that remain empty after $K$ balls thrown into $n$ urns. According to [15],
$Var[N_{0}]=\sum_{i=1}^{n}\left(e^{-Kp_{i}^{(t)}}-e^{-2Kp_{i}^{(t)}}\right)-K\left(\sum_{i=1}^{n}p_{i}^{(t)}e^{-Kp_{i}^{(t)}}\right)^{2}.$
@@
$Var[N]=Var[n-N_{0}]=Var[N_{0}]=\sum_{i=1}^{n}\left(e^{-Kp_{i}^{(t)}}-e^{-2Kp_{i}^{(t)}}\right)-K\left(\sum_{i=1}^{n}p_{i}^{(t)}e^{-Kp_{i}^{(t)}}\right)^{2}.$
@@
$N\sim\left(m\left(1-e^{-\alpha}\right),m(e^{-\alpha}-e^{-2\alpha}-\alpha e^{-2\alpha})\right).$
In summary, for given $m$ information strands, when sequencing fixed $K$ reads, the expected ratio of successfully decoded designed strands is $\frac{N}{m}$. That is, if $m$ information strands are encoded with $\left[\frac{m^{2}}{N},m\right]$ MDS code, then theoretically speaking, sequencing $K$ reads in a noiseless channel recovers the full information.
Applying the conclusions of Theorem 1 and Theorem 2, we present in Fig. 2 a comparison of the minimum sequencing coverage depth under real channels and a uniform distribution channel. Fig. 3 visualize the variance derivative function to intuitively show the probability of successfully decoding all the digital data in a single sequencing experiment, where $f(K)\triangleq e^{-K\mathbb{E}\left[p_{i}^{(t)}\right]}-e^{-2K\mathbb{E}\left[p_{i}^{(t)}\right]}-\frac{n}{K}\left(K\mathbb{E}\left[p_{i}^{(t)}\right]\right)^{2}e^{-2K\mathbb{E}\left[p_{i}^{(t)}\right]}$.
Based on the applications of the Theorem 1 and Theorem 2, we obtain the following conclusions:
@@
Figure 3: The expectation of the the coverage depth with respect to $[n,m]$ MDS code redundancy.
Figure 4: The derivative of the variance of the coverage depth. The small figure shows the variance of decoding each oligo under different PCR amplification cycles. The trend of the variance derivative reflects the probability of successfully decoding all the data information in a single sequencing experiment when the expected sequencing sample size is used as the actual sequencing sample size.
IV-B Sequencing Coverage Depth in the Noisy Channel
Theorem 3 and Theorem 4 of this section give two lower bounds on $I_{a}^{\boldsymbol{p}_{t}}(n,m)$ in noisy channels obeying a log-normal distribution.
@@ IV-B Sequencing Coverage Depth in the Noisy Channel
Theorem 3. For given $\boldsymbol{c}$, $\boldsymbol{r}$, for any $t\geq 1$, $\beta>1$, $a>1$, let $R=\frac{n}{m}$, it holds that $P\left[I_{a}^{\boldsymbol{p}_{t}}(n,m)\leq K_{1}\left(\mu^{(t)},\sigma^{(t)},a,R\right)\right]\leq e^{-\beta}\left(1+\frac{m}{n-m}\right)$, where
$K_{1}(\mu^{(t)},\sigma^{(t)},a,R)\triangleq e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\left(\log\Bigg{(}\frac{R}{R-1}\right)-\beta\Bigg{)}.$
Proof. Since $I_{1}^{\boldsymbol{p}_{t}}(n,m)<I_{a}^{\boldsymbol{p}_{t}}(n,m)$, the following proves that
$P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)\leq K_{1}\right]\leq e^{-\beta}\left(1+\frac{m}{n-m}\right).$
@@
$e^{\log\left(\frac{n}{n-m}\right)-\beta}\cdot\mathbb{E}\left[e^{-\mathbb{E}\left[p_{i}^{(t)}\right]\cdot I_{1}^{\boldsymbol{p}_{t}}(n,m)}\right]$ $=e^{\log\left(\frac{n}{n-m}\right)-\beta}\cdot\sum_{k=1}^{\infty}e^{-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$=\sum_{k=1}^{\infty}e^{\log\left(\frac{n}{n-m}\right)-\beta-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$=\sum_{k=1}^{\left\lfloor\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right\rfloor}e^{\log\left(\frac{n}{n-m}\right)-\beta-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$+\sum_{k=1+\left\lfloor\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right\rfloor}^{\infty}e^{\log\left(\frac{n}{n-m}\right)-\beta-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$\geq\sum_{k=1}^{\left\lfloor\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right\rfloor}e^{\log\left(\frac{n}{n-m}\right)-\beta-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$\geq\sum_{k=1}^{\left\lfloor\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right\rfloor}1\cdot P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$\geq P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)\leq\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right]$
$=P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)\leq e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\cdot\log\left(\frac{n}{n-m}\right)-e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\cdot\beta\right].$
From [16], the generating function of the geometric random variable $I_{1}^{\boldsymbol{p}_{t}}(n,m)$ is given by
$G_{I_{1}^{\boldsymbol{p}_{t}}(n,m)}\left(e^{-\mathbb{E}\left[p_{i}^{(t)}\right]}\right)$ $=\mathbb{E}\left[{e^{-\mathbb{E}\left[p_{i}^{(t)}\right]\cdot{I_{1}^{\boldsymbol{p}_{t}}(n,m)}}}\right]$
$=\sum_{k=0}^{\infty}P\left[I_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]{\left(e^{-\mathbb{E}\left[p_{i}^{(t)}\right]}\right)}^{k}$
@@
$P\left[I_{a}^{\boldsymbol{p}_{t}}(n,m)\leq K_{1}\left(\mu^{(t)},\sigma^{(t)},a,R\right)\right]$ $\leq e^{\log\left(\frac{n}{n-m}\right)-\beta}\cdot\mathbb{E}\left[e^{-\mathbb{E}\left[p_{i}^{(t)}\right]\cdot I_{1}^{\boldsymbol{p}_{t}}(n,m)}\right]$
@@
Theorem 3 provides a lower bound on the expected sequencing coverage depth, independent of coding redundancy. The following theorem establishes a lower bound on the expected sequencing coverage depth in relation to coding redundancy. Let $N_{K}$ represents the number of urns that contain less than $a$ balls after $K$ rounds. When using $[n,m]$ MDS code, recovering all data means successfully decoding $m$ out of $n$ strands, implying that at most $n-m$ strands cannot be successfully decoded.
Theorem 4. For given $\boldsymbol{c}$, $\boldsymbol{r}$, for any $a>1$, $t\geq 1$, let $R=\frac{n}{m}$, we have $\mathbb{E}[N_{K}]\leq n-m$, if $I_{a}^{\boldsymbol{p}_{t}}(n,m)>K_{2}\left(\mu^{(t)},\sigma^{(t)},a,R\right)$, where
$K_{2}\left(\mu^{(t)},\sigma^{(t)},a,R\right)\triangleq e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\Bigg{(}(a-1)-\log 2\log\left(\frac{R-1}{R}\right)+(a-1)\sqrt{-\frac{2\log 2}{a-1}\log\left(\frac{R-1}{R}\right)}\Bigg{)}.$
@@
Lemma 1. For $K>(a-1)\cdot\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]$, we have $\mathbb{E}\left[N_{K}\right]\leq n-m$, if $\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}e^{-\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}}\leq\frac{1}{e}\left(1-\frac{1}{R}\right)^{\frac{\log 2}{a-1}}$.
Proof. We consider this question in the context of the urn problem. According to Claim 3 in [7], the probability of the $i$-th urn containing less than $a$ balls after $K$ rounds is
$q_{i}=\sum_{j=0}^{a-1}\binom{K}{j}{p_{i}}^{j}(1-p_{i})^{K-j}\leq e^{KD\left(\frac{a-1}{K}\Big{|}\Big{|}p_{i}\right)},$
@@
where the second inequality follows from Jensen’s inequality.
Hence, $\mathbb{E}[N_{K}]\leq n-k=n(1-R)$ holds if and only if
@@
Proof of Theorem 4. Let $x=\frac{1}{e}\left(1-\frac{1}{R}\right)^{\frac{\log 2}{a-1}}$, $y=\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}$, according to Claim 7 in [7], the equation
$-\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}e^{-\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}}=-ye^{-y}=-x=-\frac{1}{e}\left(1-\frac{1}{R}\right)^{\frac{\log 2}{a-1}}$
has two solutions $y=-W_{0}(-x)$ and $y=-W_{-1}(-x)$, and Lemma 1 holds if and only if $y\geq-W_{-1}(-x)$. Then, from Claim 8 in [7], we have that,
$-W_{-1}(-x)<1+\sqrt{-\frac{2\log 2}{a-1}\log\left(1-\frac{1}{R}\right)}-\frac{\log 2}{a-1}\log\left(1-\frac{1}{R}\right).$
Hence, according to the expression of $y$, a sufficient condition for $\mathbb{E}\left[N_{K}\right]\leq n-m$ is
$\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}\geq 1+\sqrt{-\frac{2\log 2}{a-1}\log\left(1-\frac{1}{R}\right)}-\frac{\log 2}{a-1}\log\left(1-\frac{1}{R}\right),$
@@
$K\geq\frac{1}{\mathbb{E}\left[p_{i}^{(t)}\right]}\cdot(a-1)-\frac{1}{\mathbb{E}\left[p_{i}^{(t)}\right]}\cdot\log 2\log\left(1-\frac{1}{R}\right)+\frac{1}{\mathbb{E}\left[p_{i}^{(t)}\right]}\cdot(a-1)\sqrt{-\frac{2\log 2}{a-1}\log\left(1-\frac{1}{R}\right)}.$
@@
$K\geq\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot(a-1)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log 2\log\left(1-\frac{1}{R}\right)+\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot(a-1)\sqrt{-\frac{2\log 2}{a-1}\log\left(1-\frac{1}{R}\right)},$
@@
$e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\left((a-1)-\log 2\log\left(\frac{R-1}{R}\right)+(a-1)\sqrt{-\frac{2\log 2}{a-1}\log\left(\frac{R-1}{R}\right)}\right).$
@@
We will visually compare the two lower bounds of the expected sequencing coverage depth that we prove with the maximum sequencing coverage depth in the real channel obtained through 1000 Monte Carlo simulations, as shown in the fig. 5 below.
(a) R=1.25 a=2
(b) R=2 a=2
Figure 5: Visualization of the theoretical lower bounds of the expected sequencing coverage depth and 1000 Monte Carlo simulations of the maximum sequencing coverage depth in the real channel.

### v2

@@ IV-A MDS Coverage Depth Problem in the Noiseless Channel
Theorem 1. For any channel probability distribution $\boldsymbol{p}_{t}=(p_{1}^{(t)},p_{2}^{(t)},\dots,p_{n}^{(t)})$ and any $K\geq n\geq m\geq 1$, it holds that
@@
$\mathbb{E}[X_{i}^{(K)}]=1-\left(1-p_{i}^{(t)}\right)^{K}\approx 1-e^{K\cdot p_{i}^{(t)}},$
@@
$\mathbb{E}[N]=\mathbb{E}\left[\sum_{i=1}^{n}X_{i}^{(K)}\right]=\sum_{i=1}^{n}\mathbb{E}[X_{i}^{(K)}]=\sum_{i=1}^{n}\left(1-e^{K\cdot p_{i}^{(t)}}\right).$
The value of $\mathrm{Var}[N]$: Inspired by [18], we first investigate the number of urns, denote it as $N_{0}$, that remain empty after throwing $K$ balls into $n$ urns. According to [18],
$\mathrm{Var}[N_{0}]=\sum_{i=1}^{n}\left(e^{-Kp_{i}^{(t)}}-e^{-2Kp_{i}^{(t)}}\right)-K\left(\sum_{i=1}^{n}p_{i}^{(t)}e^{-Kp_{i}^{(t)}}\right)^{2}.$
@@
$\mathrm{Var}[N]=\mathrm{Var}[n-N_{0}]=\mathrm{Var}[N_{0}]=\sum_{i=1}^{n}\left(e^{-Kp_{i}^{(t)}}-e^{-2Kp_{i}^{(t)}}\right)-K\left(\sum_{i=1}^{n}p_{i}^{(t)}e^{-Kp_{i}^{(t)}}\right)^{2}.$
@@
$N\sim\left(n\left(1-e^{-\alpha}\right),n(e^{-\alpha}-e^{-2\alpha}-\alpha e^{-2\alpha})\right).$
In summary, for $n$ designed strands, when sequencing fixed $K$ reads, the expected ratio of successfully decoding designed strands is $\frac{\mathbb{E}[N]}{n}$. That is, if $m$ information strands are encoded with the $\left[\frac{n\cdot m}{\mathbb{E}[N]},m\right]$ MDS code, then theoretically speaking, sequencing $K$ reads in a noiseless channel recovers complete information.
Applying the conclusions of Theorem 1 and Theorem 2, we present in Fig. 3 a comparison of the minimum sequencing coverage depth under real channels and a uniform distribution channel. Fig. 4 visualizes the variance derivative function to intuitively show the probability of successfully decoding all digital data in a single sequencing experiment.
Based on the applications of Theorem 1 and Theorem 2, we obtain the following conclusions:
@@
Figure 3: Expected coverage depth required for successful decoding under different code rates. The Uniform dist. curve represents the relationship between coverage depth and code rate in a uniform distribution channel. The PCR10, PCR30, and PCR60 curves correspond to the empirical channels derived from Dataset PCR10, Dataset PCR30, and Dataset PCR60, respectively, showing how coverage depth varies with code rate. Three points denote Monte Carlo simulations performed under the respective channel with code rate $R=0.5$.
Figure 4: The inner plot illustrates the probability that a single experiment fails to decode all information when sequencing is performed at the expected sample size. This probability is quantified by the variance of each designed strand under that sequencing sample size, denoted as $f(K)\triangleq e^{-K\mathbb{E}\left[p_{i}^{(t)}\right]}-e^{-2K\mathbb{E}\left[p_{i}^{(t)}\right]}-\frac{n}{K}\left(K\mathbb{E}\left[p_{i}^{(t)}\right]\right)^{2}e^{-2K\mathbb{E}\left[p_{i}^{(t)}\right]}$. The outer plot depicts the trend of this probability, i.e., the derivative $f^{\prime}(K)$. The outer plot indicates that the variance has only one peak, corresponding to the maximum variance point as annotated in the figure.
IV-B MDS Coverage Depth Problem in the Noisy Channel
Theorem 3 and Theorem 4 of this section present two lower bounds on $K_{a}^{\boldsymbol{p}_{t}}(n,m)$ in noisy channels obeying a log-normal distribution.
@@ IV-B MDS Coverage Depth Problem in the Noisy Channel
Theorem 3. For given $\boldsymbol{c}$, $\boldsymbol{r}$, for any $t\geq 1$, $\beta>1$, $a>1$, let $R=\frac{m}{n}$, it holds that $P\left[K_{a}^{\boldsymbol{p}_{t}}(n,m)\leq K_{1}\left(\mu^{(t)},\sigma^{(t)},a,R\right)\right]\leq e^{-\beta}\left(1+\frac{m}{n-m}\right)$, where
$K_{1}(\mu^{(t)},\sigma^{(t)},a,R)\triangleq e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\left(\log\Bigg{(}\frac{1}{1-R}\right)-\beta\Bigg{)}.$
Proof. Since $K_{1}^{\boldsymbol{p}_{t}}(n,m)<K_{a}^{\boldsymbol{p}_{t}}(n,m)$, the following proves that
$P\left[K_{1}^{\boldsymbol{p}_{t}}(n,m)\leq K_{1}\right]\leq e^{-\beta}\left(1+\frac{m}{n-m}\right).$
@@
$e^{\log\left(\frac{n}{n-m}\right)-\beta}\cdot\mathbb{E}\left[e^{-\mathbb{E}\left[p_{i}^{(t)}\right]\cdot K_{1}^{\boldsymbol{p}_{t}}(n,m)}\right]$ $=e^{\log\left(\frac{n}{n-m}\right)-\beta}\cdot\sum_{k=1}^{\infty}e^{-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[K_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$=\sum_{k=1}^{\infty}e^{\log\left(\frac{n}{n-m}\right)-\beta-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[K_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$=\sum_{k=1}^{\left\lfloor\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right\rfloor}e^{\log\left(\frac{n}{n-m}\right)-\beta-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[K_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$+\sum_{k=1+\left\lfloor\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right\rfloor}^{\infty}e^{\log\left(\frac{n}{n-m}\right)-\beta-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[K_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$\geq\sum_{k=1}^{\left\lfloor\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right\rfloor}e^{\log\left(\frac{n}{n-m}\right)-\beta-k\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}P\left[K_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$\geq\sum_{k=1}^{\left\lfloor\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right\rfloor}1\cdot P\left[K_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]$
$\geq P\left[K_{1}^{\boldsymbol{p}_{t}}(n,m)\leq\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log\left(\frac{n}{n-m}\right)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\beta\right]$
$=P\left[K_{1}^{\boldsymbol{p}_{t}}(n,m)\leq e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\cdot\log\left(\frac{n}{n-m}\right)-e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\cdot\beta\right].$
From [20], the generating function of the geometric random variable $K_{1}^{\boldsymbol{p}_{t}}(n,m)$ is given by
$G_{K_{1}^{\boldsymbol{p}_{t}}(n,m)}\left(e^{-\mathbb{E}\left[p_{i}^{(t)}\right]}\right)$ $=\mathbb{E}\left[{e^{-\mathbb{E}\left[p_{i}^{(t)}\right]\cdot{K_{1}^{\boldsymbol{p}_{t}}(n,m)}}}\right]$
$=\sum_{k=0}^{\infty}P\left[K_{1}^{\boldsymbol{p}_{t}}(n,m)=k\right]{\left(e^{-\mathbb{E}\left[p_{i}^{(t)}\right]}\right)}^{k}$
@@
$P\left[K_{a}^{\boldsymbol{p}_{t}}(n,m)\leq K_{1}\left(\mu^{(t)},\sigma^{(t)},a,R\right)\right]$ $\leq e^{\log\left(\frac{n}{n-m}\right)-\beta}\cdot\mathbb{E}\left[e^{-\mathbb{E}\left[p_{i}^{(t)}\right]\cdot K_{1}^{\boldsymbol{p}_{t}}(n,m)}\right]$
@@
Let $N_{K}$ represent the number of urns that contain less than $a$ balls after $K$ rounds. When using an $[n,m]$ MDS code, recovering all data means successfully decoding $m$ out of $n$ strands, implying that at most $n-m$ strands cannot be successfully decoded.
Theorem 4. For given $\boldsymbol{c}$, $\boldsymbol{r}$, for any $a>1$, $t\geq 1$, let $R=\frac{m}{n}$, we have $\mathbb{E}[N_{K}]\leq n-m$, if $K_{a}^{\boldsymbol{p}_{t}}(n,m)>K_{2}\left(\mu^{(t)},\sigma^{(t)},a,R\right)$, where
$K_{2}\left(\mu^{(t)},\sigma^{(t)},a,R\right)\triangleq e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\Bigg{(}(a-1)-\log 2\log(1-R)+(a-1)\sqrt{-\frac{2\log 2}{a-1}\log(1-R)}\Bigg{)}.$
@@
Lemma 1. For $K>(a-1)\cdot\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]$, we have $\mathbb{E}\left[N_{K}\right]\leq n-m$, if $\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}e^{-\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}}\leq\frac{1}{e}\left(1-R\right)^{\frac{\log 2}{a-1}}$.
Proof. For clarity in writing, we omit the superscript of $p_{i}^{(t)}$ in the proof of Lemma 1 and denote it as $p_{i}$. We consider this question in the context of the urn problem. According to Claim 2 in [5], the probability of the $i$-th urn containing less than $a$ balls after $K$ rounds is
$q_{i}=\sum_{j=0}^{a-1}\binom{K}{j}{p_{i}}^{j}(1-p_{i})^{K-j}\leq e^{-KD\left(\frac{a-1}{K}\Big{|}\Big{|}p_{i}\right)},$
@@
let $f(p_{i})=e^{-(a-1)\log_{2}\frac{a-1}{Kp_{i}}-(K-(a-1))\log_{2}\frac{K-(a-1)}{K\left(1-p_{i}\right)}}$, since $f(p)$ is a convex function, $e^{-f(p)}$ is a log-concave function, therefore, the second inequality follows from Jensen’s inequality.
Hence, $\mathbb{E}[N_{K}]\leq n-m=n(1-R)$ holds if and only if
@@
Proof of Theorem 4. Let $x=\frac{1}{e}\left(1-R\right)^{\frac{\log 2}{a-1}}$, $y=\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}$, according to Claim 7 in [5], the equation
$-\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}e^{-\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}}=-ye^{-y}=-x=-\frac{1}{e}\left(1-R\right)^{\frac{\log 2}{a-1}}$
has two solutions $y=-W_{0}(-x)$ and $y=-W_{-1}(-x)$. From Claim 8 in [5], we have that,
$-W_{-1}(-x)<1+\sqrt{-\frac{2\log 2}{a-1}\log\left(1-R\right)}-\frac{\log 2}{a-1}\log\left(1-R\right).$
The function $-ye^{-y}$ is both continuous and monotonically increasing for $y$. Therefore, for any given $R=\frac{m}{n}$, only the branch $W_{-1}$ is applicable. This means that for $K\geq(a-1)\cdot\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]$, Lemma 1 holds if and only if $y\geq-W_{-1}(-x)$. Hence, according to the expression of $y$, a sufficient condition for $\mathbb{E}\left[N_{K}\right]\leq n-m$ is
$\frac{K\cdot\mathbb{E}\left[p_{i}^{(t)}\right]}{a-1}\geq 1+\sqrt{-\frac{2\log 2}{a-1}\log\left(1-R\right)}-\frac{\log 2}{a-1}\log\left(1-R\right),$
@@
$K\geq\frac{1}{\mathbb{E}\left[p_{i}^{(t)}\right]}\cdot(a-1)-\frac{1}{\mathbb{E}\left[p_{i}^{(t)}\right]}\cdot\log 2\log\left(1-R\right)+\frac{1}{\mathbb{E}\left[p_{i}^{(t)}\right]}\cdot(a-1)\sqrt{-\frac{2\log 2}{a-1}\log\left(1-R\right)}.$
@@
$K\geq\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot(a-1)-\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot\log 2\log\left(1-R\right)+\mathbb{E}\left[\frac{1}{p_{i}^{(t)}}\right]\cdot(a-1)\sqrt{-\frac{2\log 2}{a-1}\log\left(1-R\right)},$
@@
$e^{-\mu^{(t)}+\frac{{\sigma^{(t)}}^{2}}{2}}\left((a-1)-\log 2\log(1-R)+(a-1)\sqrt{-\frac{2\log 2}{a-1}\log(1-R)}\right).$
@@
We compare the two lower bounds of the expected sequencing coverage depth that we prove (i.e., $K_{1}$ and $K_{2}$ respectively) with the maximum sequencing coverage depth in the real channel obtained through 1000 Monte Carlo simulations (i.e., sim-max) under the condition that $a=2$, $R=0.8$, as shown in Table II below.
TABLE II: Simulation and theoretical bounds ($a=2$, $R=0.8$).
Cycles $K_{1}$ $K_{2}$ sim-max
PCR10 3.33 7.51 6.12
PCR30 4.32 9.75 7.68
PCR60 7.8 17.61 12.85